OpenStudy (babyx3boo):

(x^1/3+y^-2/3)(x^4/3-y^-1/3)

OpenStudy (babyx3boo):

@TheSmartOne

OpenStudy (rizags):

what are you asking for here?

OpenStudy (babyx3boo):

I have to perform the operation like multiply it but and i'm stuck with the whole exponent fractions. Like i foiled but i feel like you can add the middle numbers o.o

OpenStudy (rizags):

Show me what you have so far

OpenStudy (rizags):

ok now show what you have do far with the expanding

OpenStudy (rizags):

thats a correct expansion, i mean you cant add those middle terms in their current state, so I'd say thats fully correct

OpenStudy (babyx3boo):

:o how come you can't add em

OpenStudy (rizags):

becuz they have different exponents

OpenStudy (rizags):

you cant just add x^2 to x, for example

OpenStudy (babyx3boo):

but don't they have the same denominator in the exponents like when you multiply the first term.

OpenStudy (rizags):

well, yes, they do, but you still cannot add them unless the exponents are exactly the same, like you can add x^2 and x, or y^1/3 and y^2/3

OpenStudy (rizags):

*can't

OpenStudy (babyx3boo):

Oh alright lol and i have at least two other problems :o mind helping?

OpenStudy (rizags):

sure i'll help! fan plz if you want

OpenStudy (babyx3boo):

Of course!

OpenStudy (babyx3boo):

Okay the question is m^2/3/m^-8/5 my answer was m^-6 but on the board it said m^2 :o

OpenStudy (rizags):

is this the problem: \[\frac{m^{\frac{2}{3}}}{m^{\frac{-8}{5}}}\]

OpenStudy (babyx3boo):

Yessss

OpenStudy (babyx3boo):

wait is it since we have to divide you got to subtract therefore you gotta change the sign o.o?

OpenStudy (rizags):

well, it becomes \[\huge\frac{ 2 }{ 3 }-\frac{ -8 }{ 5 }\]

OpenStudy (rizags):

or, \[\large\frac{2}{3}+\frac{8}{5}\]

OpenStudy (babyx3boo):

oh wait my bad the first fraction is 2/5

OpenStudy (babyx3boo):

not 2/3

OpenStudy (rizags):

ok, so its just \[\huge m^{\frac{2}{5}+\frac{8}{5}}\] can you solve that?

OpenStudy (babyx3boo):

yes 10/5 reducing to 2 which equals m^2 :)

OpenStudy (rizags):

riiiiiiiite

OpenStudy (babyx3boo):

Thank you now the other question i don't know how to start

OpenStudy (rizags):

um, this isn't really a "solvable" thing so to speak.

OpenStudy (babyx3boo):

what you mean?

OpenStudy (rizags):

like, what exactly are you supposed to do with this? simplify?

OpenStudy (babyx3boo):

I don't know it says perform the operation it's in the same section as the ones we did

OpenStudy (rizags):

i mean, i suppose you could rewrite it like this: \[\huge \frac{\frac{1}{a^2}+\frac{b}{a}-b^2}{\frac{a^2}{b}}\]

OpenStudy (babyx3boo):

that's kinda confusing o.o

OpenStudy (rizags):

ur rite xD ummmmm what class is this for?

OpenStudy (babyx3boo):

Algebra 2 lol

OpenStudy (rizags):

i'm also in algebra 2 xD but this does not look simplifiable to me

OpenStudy (babyx3boo):

It's a test review and i'm like so lost in that one but thank you so much for your help :) if i have another question can i tag u ?

OpenStudy (rizags):

yes ofc!

OpenStudy (babyx3boo):

Thank you :)

OpenStudy (rizags):

np :)